Minutes

Meeting Date File
May 22, 2015

137.  Switching back from Argon ions to Lead ions in the injector complex, SU activities in the SPS north area during LS1 and YETS 2014-2015, Consolidation plan for the SPS electrical distribution, Decabling status for the PSB, PS and SPS, T2 & T6 movement in TCC2, AD CCC cool-down status, New PS booster orbit corrector converters - LIU feedback.

May 8, 2015

136. SPSC Report, AWAKE Technical status report, AWAKE budget review, responsibility for DC cables and busbars, status of COLDEX reinstallation, Additional PSB orbit corrector power converters, new frequencies for 4G network underground, progress of the AD CCC, T2 position in TCC2.

April 24, 2015

135. Mobile network switch debrief, AD Cryogenic Current Comparator status, Decabling Status for the PSB, PS and SPS, Status of the Marguerite System of the East Area, New access right conditions for the SPS and PS Complex.

April 17, 2015

134. List of de-cabling responsibles for equipment group, Status of the leaking 80 MHz PS cavity, Machine interlock for COLDEX, n_TOF target ventilation and access system, photographs of COLDEX intervention.

April 10, 2015

133. Update on the 80 MHz cavity in PS sector 8, Update on the CCC router redundancy, Realignment of COLDEX's by-pass and required corrective interventions, summary of successful completion of NA61/SHINE Argon run and plans for the repair of the VTX-2 helium pump, statement on de-cabling campaign in the injector complex.

March 20, 2015

132. Physics case for 2 GeV beams at ISOLDE, 2 GeV transfer line from PSB to ISOLDE, ISOLDE targets and dumps for 2 GeV beams.

March 13, 2015

131. Commissioning plans for HIE-Isolde, Recommissioning of REX-ISOLDE, CE updates on BA1 and BA2 Consolidation, length of YETS2015/2016.

March 6, 2015

130. PS/PS Booster wire scanner status report, INTC report ISOLDE, INTC Report n_TOF, reference measurement of PS magnets to validate radial shift, TPST  collimator exchange in SPS, AD BCCCA installation and discussion, Status of fiche réflexe, RP storage and waste capacity.

February 20, 2015

129. Status of the POPS capacitors, EATM report, CCC IT routers redundancy, SPS rotating collimator aperture for slow extraction, Mandate for the SPS fire safety WG.

February 13, 2015

128. 2016 AD shutdown work in preparation for ELENA, Upgrade of the cooling power in the SPS, First deliverz of argon at 150 AGeV/c.

February 6, 2015

127. Repair of the main Meyrin sub-station after the event of Oct 16th, LINAC4 source status, LINAC4 12 MeV status, IMPACT during run 2.

January 30, 2015

126. Status report on SPS rotating collimator test during 2015, SPSC report.

January 23, 2015

125. SPS doublet beam compatibility with users, ECRs, 80 MHz PS SS88 Cavity Issue, ELENA/g-bar cooling system update and needs from EN-CV.

January 9, 2015

124. LIU-SPS Project Updates, Short and medium plan to prevent flood in BAs and monitoring for CERN tunnels, North Area primary ion safety system global validation tests, report on the BE-CO smooth upgrades campaign 5-6-7 January

December 12, 2014

123. CERN mobile network migration, Working Group on Recommissioning of the Injector Complex after an extended stop, update on the SPS scrubbing run.

December 5, 2014

122. AD CCC Update, Impact of other projects of the CCC Cryostat manufacturing schedule, TE-VSC position on the AD sectorization for the CCC, FGC power converter control issues, AWAKE status, Research Board Outcome, Accelerator Schedule for 2015.

November 28, 2014

121. A cryogenic current comparator for AD, report on the outcome of the ELENA CSR, ELENA project technical review summary, COLDEX/Crab cavity arbitration in SPS, cost estimate for extended north area run in 2015.

November 21, 2014

120. MEDICIS update, Report from the FOM, The CERN Neutrino Platform, EHN1 Extension for neutrino detectors R&D experiments, Cryostats and Cryogenics for large liquid argon-based neutrino experiments, TIDVG 2005 endoscopy.

November 14, 2014

119. SHIP: Short presentation of the proposed experiment, planning and next steps, SHIP: impact on beam line design, extraction and SPS Proton budget, SHIP Target station and experimental facility, SHIP Civil Engineering constraints and critical planning for LS2, Preliminary results for a-C coating of SPS magnet vacuum chambers, preliminary results from COLDEX following the scrubbing run, Mandate for the working group on re-commissioning of the injector complex after a prolonged period of maintenance.

November 7, 2014

118. Report of the HIE-ISOLDE Cost and Schedule Review, FOM Report, SPSC Report, INTC: ISOLDE & n_TOF Reports

Pages

You are here